# Gradient Estimation with Discrete Stein Operators

### v7

@@ 1 Introduction
Modern machine learning relies heavily on gradient methods to optimize the parameters of a learning system. However, exact gradient computation is often difficult. For example, in variational inference for training latent variable models (Paisley et al., 2012; Kingma and Welling, 2014), policy gradient algorithms in reinforcement learning (Williams, 1992), and combinatorial optimization (Niepert et al., 2021), the exact gradient features an intractable sum or integral introduced by an expectation under an evolving probability distribution. To make progress, one resorts to estimating the gradient by drawing samples from that distribution (Schulman et al., 2015; Mohamed et al., 2020).
The two main classes of gradient estimators used in machine learning are the pathwise or reparameterization gradient estimators (Kingma and Welling, 2014; Rezende et al., 2014; Titsias and Lázaro-Gredilla, 2014) and the REINFORCE or score function estimators (Williams, 1992; Glynn, 1990). The pathwise estimators have shown great success in training variational autoencoders (Kingma and Welling, 2014) but are only applicable to continuous probability distributions. The REINFORCE estimators are more general-purpose and easily accommodate discrete distributions but often suffer from excessively high variance.
To improve the quality of REINFORCE estimators, we develop a new variance reduction technique for discrete expectations. Our method is based upon Stein operators (see, e.g., Stein, 1972; Anastasiou et al., 2021), computable functionals that generate mean-zero functions under a target distribution and provide a natural way of designing control variates (CVs) for stochastic estimation.
We first provide a general recipe for constructing practical Stein operators for discrete distributions (Table 1), generalizing the prior literature (Brown and Xia, 2001; Holmes, 2004; Eichelsbacher and Reinert, 2008; Bresler and Nagaraj, 2019; Reinert and Ross, 2019; Yang et al., 2018; Barp et al., 2019). We then develop a gradient estimation framework—RODEO—that augments REINFORCE estimators with mean-zero CVs generated from Stein operators. Finally, inspired by Double CV (Titsias and Shi, 2022), we extend our method to develop CVs for REINFORCE leave-one-out estimators (Salimans and Knowles, 2014; Kool et al., 2019) to further reduce the variance.
@@
We evaluate RODEO on 15 benchmark tasks, including training binary variational autoencoders (VAEs) with one or more stochastic layers. In most cases and with the same number of function evaluations, RODEO delivers lower variance and better training objectives than the state-of-the-art gradient estimators DisARM (Dong et al., 2020; Yin et al., 2020), ARMS (Dimitriev and Zhou, 2021), Double CV (Titsias and Shi, 2022), and RELAX (Grathwohl et al., 2018).
@@ 2 Background
We consider the problem of maximizing the objective function $\mathbb{E}_{q_{\eta}}[f(x)]$ with respect to the parameters $\eta$ of a discrete distribution $q_{\eta}(x)$. Throughout the paper we assume $f(x)$ is a differentiable function of real-valued inputs $x\in\mathbb{R}^{d}$ but is only evaluated at a discrete subset $\mathcal{X}^{d}$ due to the discreteness of $q_{\eta}$. [^1] This assumption holds for many discrete probabilistic models (Grathwohl et al., 2021) including the binary VAEs of Section 6. Exact computation of the expectation is typically intractable due to the complex nature of $f(x)$ and $q_{\eta}(x)$. The standard workaround is to rewrite the gradient as $\nabla_{\eta}\mathbb{E}_{q_{\eta}}[f(x)]=\mathbb{E}_{q_{\eta}}[f(x)\nabla_{\eta}\log q_{\eta}(x)]$ and employ the Monte Carlo gradient estimator known as the score function or REINFORCE estimator (Glynn, 1990; Williams, 1992):
@@
Here, $b$ is a constant called the “baseline” introduced to reduce the variance of the estimator by reducing the scaling effect of $f(x^{(k)})$. Since $\mathbb{E}_{q_{\eta}}[\nabla_{\eta}\log q_{\eta}(x)]=0$, the REINFORCE estimator is unbiased for any choice of $b$, and the term $b\nabla_{\eta}\log q_{\eta}(x)$ is known as a control variate (CV) (Owen, 2013, Ch. 8). The optimal baseline can be estimated using additional function evaluations (Weaver and Tao, 2001; Bengio et al., 2013; Ranganath et al., 2014). A simpler approach is to use moving averages of $f$ from historical evaluations or to train function approximators to mimic those values (Mnih and Gregor, 2014). When $K\geq 2$, a powerful variant of REINFORCE is obtained by replacing $b$ with the leave-one-out average of function values:
@@
This approach is often called the REINFORCE leave-one-out (RLOO) estimator (Salimans and Knowles, 2014; Kool et al., 2019; Richter et al., 2020) and was recently observed to have very strong performance in training discrete latent variable models (Richter et al., 2020; Dong et al., 2020).
All the above methods construct a baseline that is independent of the point $x^{(k)}$ under consideration, but there are other ways to preserve the unbiasedness of the estimator. We are free to use a sample-dependent baseline $b(x^{(k)})$ as long as the expectation $c\triangleq\mathbb{E}_{q_{\eta}}[b(x^{(k)})\nabla_{\eta}\log q_{\eta}(x^{(k)})]$ is easily computed or has a low-variance unbiased estimate. In this case, we can correct for any bias introduced by $b(x^{(k)})$ by adding in this expectation: $\frac{1}{K}\sum_{k=1}^{K}(f(x^{(k)})-b(x^{(k)}))\nabla_{\eta}\log q_{\eta}(x^{(k)})+c.$ For example, $b(x^{(k)})$ can be a lower bound or a Taylor expansion of $f(x^{(k)})$ (Paisley et al., 2012; Gu et al., 2016). Taking this a step further, the Double CV estimator (Titsias and Shi, 2022) proposed to treat $f(x^{(k)})-b(x^{(k)})$ as the effective objective function and apply the leave-one-out idea:
@@
The resulting estimator adds two CVs to RLOO: the global CV $b_{k}(x^{(k)})\nabla_{\eta}\log q_{\eta}(x^{(k)})$ and the local CV $b_{k}(x^{(j)})$. Intuitively, $b_{k}(x^{(j)})$ is aimed at reducing the variance of the LOO average. This is motivated by the fact that replacing $\frac{1}{K-1}\sum_{j\neq k}f(x^{(j)})$ with $\mathbb{E}_{q_{\eta}}[f]$ in RLOO leads to lower variance (Titsias and Shi, 2022, Prop 1). To obtain a tractable correction term $c$, Titsias and Shi (2022) adopt a linear design of $b_{k}$: $b_{k}(x)=\alpha\cdot\frac{1}{K-1}\sum_{j\neq k}f(x^{(}j))(x-\mu)$ for $\mu=\mathbb{E}_{q_{\eta}}[x]$ and a coefficient $\alpha$.
@@
In the following two sections, we will introduce a method that allows us to use very flexible CVs while still maintaining a tractable correction term. Our method enables online adaptation of CVs to minimize gradient variance (similar to RELAX (Grathwohl et al., 2018)) but does not assume $q_{\eta}$ has a continuous reparameterization. We then apply it to generalize the linear CVs in Double CV to very flexible ones such as neural networks. Moreover, we provide an effective CV design based on surrogate functions that requires no additional evaluation of $f$ compared to RLOO.
@@ 3 Control Variates from Discrete Stein Operators
for any integrable function $h$. We overload our notation so that, for any suitable matrix $A$, $(Ah)(x)\triangleq\sum_{y}A_{xy}h(y)$. In other words, for any integrable $h$, the function $(P-I)h$ is a valid CV as it has known mean under $q$. Moreover, the linear operator $P-I$ is an example of a Stein operator (Stein, 1972; Anastasiou et al., 2021) in the sense that it generates mean-zero functions under $q$. In fact, both Stein et al. (2004) and Dellaportas and Kontoyiannis (2012) developed CVs of the form $(P-I)h$ based on reversible discrete-time Markov chains and linear input functions $h(x)=x_{i}$.
More generally, Barbour (1988) and Henderson (1997) observed that if we identify a continuous-time Markov chain $(X^{(t)})_{t\geq 0}$ with $q$ as its stationary distribution, then the generator $A$, defined via
@@
satisfies $A^{\top}q=0$ and hence $\mathbb{E}_{q}[Ah]=0$ for all integrable $h$. Therefore, $A$ is also a Stein operator suitable for generating CVs. Moreover, since any discrete-time chain with transition matrix $P$ can be embedded into a continuous-time chain with transition rate matrix $A=P-I$, this continuous-time construction is strictly more general (Stirzaker, 2005, Ch. 4).
@@ 3.1 Discrete Stein operators
Gibbs Stein operator   The transition kernel of the random-scan Gibbs sampler with stationary distribution $q$ (see, e.g., Geyer, 2011) is
@@
In the binary variable setting, 4 recovers the operator Bresler and Nagaraj (2019); Reinert and Ross (2019) used to bound distances between the stationary distributions of Glauber dynamics Markov chains.
Zanella Stein operator   Zanella (2020) studied continuous-time Markov chains with generator
@@
where $\mathcal{N}_{x}$ denotes the transition neighborhood of $x$ and $\kappa$ is a continuous positive function satisfying $\kappa(t)=t\kappa(1/t)$. Conveniently, the neighborhood structure can be arbitrarily sparse. Moreover, the $\kappa$ constraint ensures that the chain satisfies detailed balance (i.e., $q(x)A_{xy}=q(y)A_{yx}$) and hence that $A^{\top}q=0$. Hodgkinson et al. (2020) call the associated operator the Zanella Stein operator:
@@
Important special cases include the minimum probability flow (Sohl-Dickstein et al., 2011) (MPF) Stein operator ($\kappa(t)=\sqrt{t}$) discussed in Barp et al. (2019) and the Barker (Barker, 1965) Stein operator ($\kappa(t)=\frac{t}{t+1}$).
@@
Then the product birth-death process (Karlin and McGregor, 1957) on $\mathcal{X}^{d}$ with birth rates $b_{i,x}=\frac{q(\texttt{inc}_{i}({x}))}{q(x)}$, death rates $d_{i,x}=1$, and generator
@@
An analogous operator is available for countably infinite $\mathcal{X}$ (Hodgkinson et al., 2020), and Brown and Xia (2001); Holmes (2004); Eichelsbacher and Reinert (2008) used related operators to characterize convergence to discrete target distributions. Moreover, by substituting $h(x)=g(x)-g(\texttt{inc}({x}))$ in 7, we recover the difference Stein operator proposed by Yang et al. (2018) without reference to birth-death processes:
@@ 4 Gradient Estimation with Discrete Stein Operators
where we now only need to learn a scalar-valued function $h$. Notably, when $h$ exactly equals $f$ and $A=P-I$ for any discrete time Markov chain kernel $P$, our CV adjustment amounts to Rao-Blackwellization (Owen, 2013, Sec. 8.7), as we end up replacing $f(x)\nabla_{\eta_{i}}\log q_{\eta}(x)$ with its conditional expectation $P(f\nabla_{\eta_{i}}\log q_{\eta})(x)=\mathbb{E}_{X_{t+1}|X_{t}=x}[f(X_{t+1})\nabla_{\eta_{i}}\log q_{\eta}(X_{t+1})]$. This yields a guaranteed variance reduction.
Surrogate function design   Based on the above reasoning, we can view $h$ as a surrogate for $f$. We avoid directly setting $h=f$ because our Stein operators evaluate $h$ at all neighbors of the sample points, and $f$ can be very expensive to evaluate (see, e.g., Shazeer et al., 2017). To avoid this problem, we first observe that $\tilde{h}$ 11 can be made sample-specific, i.e., we can use a different $\tilde{h}_{k}$ for each sample point $x^{(k)}$:
@@
Here $H$ belongs to a flexible parametric family of functions such as neural networks and is chosen to have significantly lower cost than $f$. $y$ will take on the values of $x^{(k)}$ and its neighbors in the Markov chain. We omit $x^{(k)}$ in the sum 13 to ensure that the function $h_{k}$ is independent of $x^{(k)}$ and hence that $\mathbb{E}_{q_{\eta}}[\frac{1}{K}\sum_{k=1}^{K}(A\tilde{h}_{k})(x^{(k)})]=0$. Moreover, this surrogate function design introduces no additional evaluations of $f$ beyond those required for the usual RLOO estimator. Also, as observed by Titsias and Shi (2022), for many applications, including VAE training, where $f$ has parameters $\theta$ learned through gradient-based optimization, $\{\nabla f(x^{(k)})\}_{k=1}^{K}$ can be obtained “for free” from the same backpropagation used to compute $\nabla_{\theta}f_{\theta}(x^{(k)})$ (see Algorithm 1).
RODEO  We can further improve the estimator by leveraging discrete Stein operators and the above surrogate function design to construct both the global and local CVs in the Double CV framework (Titsias and Shi, 2022) (see Section 2). We call our final estimator RODEO for RLOO with Discrete StEin Operators:
@@
Optimization with RODEO   In practice, we let the functions $H$ and $H^{\star}$ share a neural network architecture with two output units. Since the estimator is unbiased, we can optimize the parameters $\gamma$ of the network in an online fashion to minimize the variance of the estimator (similar to (Grathwohl et al., 2018)). Specifically, if we denote the RODEO gradient estimate by $g_{\gamma}(x^{(1:K)})$, then $\nabla_{\gamma}\mathrm{Tr}(\mathrm{Var}(g_{\gamma}(x^{(1:K)})))=\mathbb{E}[\nabla_{\gamma}\|g_{\gamma}(x^{(1:K)})\|_{2}^{2}].$ In Algorithm 1, we use an unbiased estimate of this hyperparameter gradient, $\nabla_{\gamma}\|g_{\gamma}(x^{(1:K)})\|_{2}^{2}$, to update $\gamma$ after each optimization step of $\eta$.

### v8

@@ 1 Introduction
Modern machine learning relies heavily on gradient methods to optimize the parameters of a learning system. However, exact gradient computation is often difficult. For example, in variational inference for training latent variable models [43, 29], policy gradient algorithms in reinforcement learning [65], and combinatorial optimization [40], the exact gradient features an intractable sum or integral introduced by an expectation under an evolving probability distribution. To make progress, one resorts to estimating the gradient by drawing samples from that distribution [50, 39].
The two main classes of gradient estimators used in machine learning are the pathwise or reparameterization gradient estimators [29, 47, 58] and the REINFORCE or score function estimators [65, 18]. The pathwise estimators have shown great success in training variational autoencoders [29] but are only applicable to continuous probability distributions. The REINFORCE estimators are more general-purpose and easily accommodate discrete distributions but often suffer from excessively high variance.
To improve the quality of REINFORCE estimators, we develop a new variance reduction technique for discrete expectations. Our method is based upon Stein operators [see, e.g., 55, 1], computable functionals that generate mean-zero functions under a target distribution and provide a natural way of designing control variates (CVs) for stochastic estimation.
We first provide a general recipe for constructing practical Stein operators for discrete distributions (Table 1), generalizing the prior literature [9, 25, 16, 8, 46, 67, 6]. We then develop a gradient estimation framework—RODEO—that augments REINFORCE estimators with mean-zero CVs generated from Stein operators. Finally, inspired by Double CV [60], we extend our method to develop CVs for REINFORCE leave-one-out estimators [49, 30] to further reduce the variance.
@@
We evaluate RODEO on 15 benchmark tasks, including training binary variational autoencoders (VAEs) with one or more stochastic layers. In most cases and with the same number of function evaluations, RODEO delivers lower variance and better training objectives than the state-of-the-art gradient estimators DisARM [14, 69], ARMS [13], Double CV [60], and RELAX [20].
@@ 2 Background
We consider the problem of maximizing the objective function $\mathbb{E}_{q_{\eta}}[f(x)]$ with respect to the parameters $\eta$ of a discrete distribution $q_{\eta}(x)$. Throughout the paper we assume $f(x)$ is a differentiable function of real-valued inputs $x\in\mathbb{R}^{d}$ but is only evaluated at a discrete subset $\mathcal{X}^{d}$ due to the discreteness of $q_{\eta}$. [^1] This assumption holds for many discrete probabilistic models [21] including the binary VAEs of Section 6. Exact computation of the expectation is typically intractable due to the complex nature of $f(x)$ and $q_{\eta}(x)$. The standard workaround is to rewrite the gradient as $\nabla_{\eta}\mathbb{E}_{q_{\eta}}[f(x)]=\mathbb{E}_{q_{\eta}}[f(x)\nabla_{\eta}\log q_{\eta}(x)]$ and employ the Monte Carlo gradient estimator known as the score function or REINFORCE estimator [18, 65]:
@@
Here, $b$ is a constant called the “baseline” introduced to reduce the variance of the estimator by reducing the scaling effect of $f(x^{(k)})$. Since $\mathbb{E}_{q_{\eta}}[\nabla_{\eta}\log q_{\eta}(x)]=0$, the REINFORCE estimator is unbiased for any choice of $b$, and the term $b\nabla_{\eta}\log q_{\eta}(x)$ is known as a control variate (CV) [42, Ch. 8]. The optimal baseline can be estimated using additional function evaluations [64, 7, 45]. A simpler approach is to use moving averages of $f$ from historical evaluations or to train function approximators to mimic those values [37]. When $K\geq 2$, a powerful variant of REINFORCE is obtained by replacing $b$ with the leave-one-out average of function values:
@@
This approach is often called the REINFORCE leave-one-out (RLOO) estimator [49, 30, 48] and was recently observed to have very strong performance in training discrete latent variable models [48, 14].
All the above methods construct a baseline that is independent of the point $x^{(k)}$ under consideration, but there are other ways to preserve the unbiasedness of the estimator. We are free to use a sample-dependent baseline $b(x^{(k)})$ as long as the expectation $c\triangleq\mathbb{E}_{q_{\eta}}[b(x^{(k)})\nabla_{\eta}\log q_{\eta}(x^{(k)})]$ is easily computed or has a low-variance unbiased estimate. In this case, we can correct for any bias introduced by $b(x^{(k)})$ by adding in this expectation: $\frac{1}{K}\sum_{k=1}^{K}(f(x^{(k)})-b(x^{(k)}))\nabla_{\eta}\log q_{\eta}(x^{(k)})+c.$ For example, $b(x^{(k)})$ can be a lower bound or a Taylor expansion of $f(x^{(k)})$ [43, 22]. Taking this a step further, the Double CV estimator [60] proposed to treat $f(x^{(k)})-b(x^{(k)})$ as the effective objective function and apply the leave-one-out idea:
@@
The resulting estimator adds two CVs to RLOO: the global CV $b_{k}(x^{(k)})\nabla_{\eta}\log q_{\eta}(x^{(k)})$ and the local CV $b_{k}(x^{(j)})$. Intuitively, $b_{k}(x^{(j)})$ is aimed at reducing the variance of the LOO average. This is motivated by the fact that replacing $\frac{1}{K-1}\sum_{j\neq k}f(x^{(j)})$ with $\mathbb{E}_{q_{\eta}}[f]$ in RLOO leads to lower variance [60, Prop 1]. To obtain a tractable correction term $c$, Titsias and Shi [60] adopt a linear design of $b_{k}$: $b_{k}(x)=\alpha\cdot\frac{1}{K-1}\sum_{j\neq k}f(x^{(j)})(x-\mu)$ for $\mu=\mathbb{E}_{q_{\eta}}[x]$ and a coefficient $\alpha$.
@@
In the following two sections, we will introduce a method that allows us to use very flexible CVs while still maintaining a tractable correction term. Our method enables online adaptation of CVs to minimize gradient variance (similar to RELAX [20]) but does not assume $q_{\eta}$ has a continuous reparameterization. We then apply it to generalize the linear CVs in Double CV to very flexible ones such as neural networks. Moreover, we provide an effective CV design based on surrogate functions that requires no additional evaluation of $f$ compared to RLOO.
@@ 3 Control Variates from Discrete Stein Operators
for any integrable function $h$. We overload our notation so that, for any suitable matrix $A$, $(Ah)(x)\triangleq\sum_{y}A_{xy}h(y)$. In other words, for any integrable $h$, the function $(P-I)h$ is a valid CV as it has known mean under $q$. Moreover, the linear operator $P-I$ is an example of a Stein operator [55, 1] in the sense that it generates mean-zero functions under $q$. In fact, both Stein et al. [56] and Dellaportas and Kontoyiannis [12] developed CVs of the form $(P-I)h$ based on reversible discrete-time Markov chains and linear input functions $h(x)=x_{i}$.
More generally, Barbour [3] and Henderson [23] observed that if we identify a continuous-time Markov chain $(X^{(t)})_{t\geq 0}$ with $q$ as its stationary distribution, then the generator $A$, defined via
@@
satisfies $A^{\top}q=0$ and hence $\mathbb{E}_{q}[Ah]=0$ for all integrable $h$. Therefore, $A$ is also a Stein operator suitable for generating CVs. Moreover, since any discrete-time chain with transition matrix $P$ can be embedded into a continuous-time chain with transition rate matrix $A=P-I$, this continuous-time construction is strictly more general [57, Ch. 4].
@@ 3.1 Discrete Stein operators
Gibbs Stein operator   The transition kernel of the random-scan Gibbs sampler with stationary distribution $q$ [see, e.g., 17] is
@@
In the binary variable setting, 4 recovers the operator Bresler and Nagaraj [8], Reinert and Ross [46] used to bound distances between the stationary distributions of Glauber dynamics Markov chains.
Zanella Stein operator   Zanella [70] studied continuous-time Markov chains with generator
@@
where $\mathcal{N}_{x}$ denotes the transition neighborhood of $x$ and $\kappa$ is a continuous positive function satisfying $\kappa(t)=t\kappa(1/t)$. Conveniently, the neighborhood structure can be arbitrarily sparse. Moreover, the $\kappa$ constraint ensures that the chain satisfies detailed balance (i.e., $q(x)A_{xy}=q(y)A_{yx}$) and hence that $A^{\top}q=0$. Hodgkinson et al. [24] call the associated operator the Zanella Stein operator:
@@
Important special cases include the minimum probability flow [53] (MPF) Stein operator ($\kappa(t)=\sqrt{t}$) discussed in Barp et al. [6] and the Barker [4] Stein operator ($\kappa(t)=\frac{t}{t+1}$).
@@
Then the product birth-death process [28] on $\mathcal{X}^{d}$ with birth rates $b_{i,x}=\frac{q(\texttt{inc}_{i}({x}))}{q(x)}$, death rates $d_{i,x}=1$, and generator
@@
An analogous operator is available for countably infinite $\mathcal{X}$ [24], and Brown and Xia [9], Holmes [25], Eichelsbacher and Reinert [16] used related operators to characterize convergence to discrete target distributions. Moreover, by substituting $h(x)=g(x)-g(\texttt{inc}({x}))$ in 7, we recover the difference Stein operator proposed by Yang et al. [67] without reference to birth-death processes:
@@ 4 Gradient Estimation with Discrete Stein Operators
where we now only need to learn a scalar-valued function $h$. Notably, when $h$ exactly equals $f$ and $A=P-I$ for any discrete time Markov chain kernel $P$, our CV adjustment amounts to Rao-Blackwellization [42, Sec. 8.7], as we end up replacing $f(x)\nabla_{\eta_{i}}\log q_{\eta}(x)$ with its conditional expectation $P(f\nabla_{\eta_{i}}\log q_{\eta})(x)=\mathbb{E}_{X_{t+1}|X_{t}=x}[f(X_{t+1})\nabla_{\eta_{i}}\log q_{\eta}(X_{t+1})]$. This yields a guaranteed variance reduction.
Surrogate function design   Based on the above reasoning, we can view $h$ as a surrogate for $f$. We avoid directly setting $h=f$ because our Stein operators evaluate $h$ at all neighbors of the sample points, and $f$ can be very expensive to evaluate [see, e.g., 51]. To avoid this problem, we first observe that $\tilde{h}$ 11 can be made sample-specific, i.e., we can use a different $\tilde{h}_{k}$ for each sample point $x^{(k)}$:
@@
Here $H$ belongs to a flexible parametric family of functions such as neural networks and is chosen to have significantly lower cost than $f$. $y$ will take on the values of $x^{(k)}$ and its neighbors in the Markov chain. We omit $x^{(k)}$ in the sum 13 to ensure that the function $h_{k}$ is independent of $x^{(k)}$ and hence that $\mathbb{E}_{q_{\eta}}[\frac{1}{K}\sum_{k=1}^{K}(A\tilde{h}_{k})(x^{(k)})]=0$. Moreover, this surrogate function design introduces no additional evaluations of $f$ beyond those required for the usual RLOO estimator. Also, as observed by Titsias and Shi [60], for many applications, including VAE training, where $f$ has parameters $\theta$ learned through gradient-based optimization, $\{\nabla f(x^{(k)})\}_{k=1}^{K}$ can be obtained “for free” from the same backpropagation used to compute $\nabla_{\theta}f_{\theta}(x^{(k)})$ (see Algorithm 1).
RODEO  We can further improve the estimator by leveraging discrete Stein operators and the above surrogate function design to construct both the global and local CVs in the Double CV framework [60] (see Section 2). We call our final estimator RODEO for RLOO with Discrete StEin Operators:
@@
Optimization with RODEO   In practice, we let the functions $H$ and $H^{\star}$ share a neural network architecture with two output units. Since the estimator is unbiased, we can optimize the parameters $\gamma$ of the network in an online fashion to minimize the variance of the estimator (similar to [20]). Specifically, if we denote the RODEO gradient estimate by $g_{\gamma}(x^{(1:K)})$, then $\nabla_{\gamma}\mathrm{Tr}(\mathrm{Var}(g_{\gamma}(x^{(1:K)})))=\mathbb{E}[\nabla_{\gamma}\|g_{\gamma}(x^{(1:K)})\|_{2}^{2}].$ In Algorithm 1, we use an unbiased estimate of this hyperparameter gradient, $\nabla_{\gamma}\|g_{\gamma}(x^{(1:K)})\|_{2}^{2}$, to update $\gamma$ after each optimization step of $\eta$.